OpenStudy (anonymous):

i started out with 13x-11y=-12 what did i do wrong? -11y=-12?

OpenStudy (amistre64):

you dropped the 13x off somewhere and unbalanced your equation

OpenStudy (amistre64):

13x -11y = -12 ; -13x to both sides to keep it equal

OpenStudy (amistre64):

13x-13x -11y = -12 - 13x

OpenStudy (amistre64):

-11y = -13x -12

OpenStudy (anonymous):

then i got -11y=-25

OpenStudy (amistre64):

-13x does not comdine with -12....they are two seperate entities that dont mix

OpenStudy (amistre64):

if you have 13 apples and 12 oranges; how many apples do you have?

OpenStudy (anonymous):

13 apples so i divide -11 to both sides to get y alone?

OpenStudy (amistre64):

yes... (-11/-11)y = -13x/-11 -12/-11 y = (13/11)x + (12/11)

OpenStudy (anonymous):

the / sign means divide right?

OpenStudy (anonymous):

and you get some random mixed up number?

OpenStudy (amistre64):

yes... '/' means divide and you get a decimaled number if you want to go that route